### v1

@@ 2 Setup
Unlike Equation (2.3), our proposed fit (2.5) does not yield an estimator of the average treatment effect (2.1). This is implicit in Theorem 1 below, and was already shown by Calonico et al. (2019). There are two potential remedies to this, neither of which can be recommended in practice. First, the covariates can be demeaned on each side of the cutoff. This would follow the logic for linear regression adjustments in randomized experiments, but in this case the demeaning would amount to Nadaraya–Watson estimation at a boundary point with $\mathbf{W}_{i}$ as the outcome, thereby hampering the entire RD estimation due to the presence of smoothing bias. Second, one could average the fitted CATEs, but this again essentially amounts to Nadaraya–Watson estimation at a boundary point, now using $\widehat{\kappa}(\mathbf{w})$ as the outcome. In both cases, the convergence rate will be slow because the misspecification bias is of order $h$, and is thus impractical for inference (e.g., severe undersmoothing would be needed). Therefore, in applications, researchers should only use our proposed methods for heterogeneity analysis and rely on the established methods underlying (2.2) or (2.3) to estimate the average RD treatment effect.
Finally, a related motivation for the specification (2.7) is to control for panel or group structures. It is common in some RD applications to include fixed effects for individuals, groups, or time periods. Identification, estimation, and inference then follow from the same arguments presented in the upcoming sections. In this way, our proposed RD methodology, although developed for the cross-sectional setting, can be easily applied to panel data, provided weighted linear least squares methods with additive or interactive fixed effect are used.
@@ 3 Main Results
We turn to the theoretical properties of the regression fit in (2.5), and particularly the CATE estimator (2.6). We provide a complete set of practical tools for identification, estimation, and inference. The following standard regularity conditions are imposed; recall that $c=0$ without loss of generality, and let $x_{\mathtt{L}}<0<x_{\mathtt{U}}$. All omitted details are given in the supplemental appendix (see Section SA1.3).
@@
The data $(Y_{i},X_{i},\mathbf{W}_{i}^{\prime}),i=1,\ldots,n$, where $T_{i}=\mathbbm{1}\{X_{i}\geq 0\}$, $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$, and the $\mathbf{W}_{i}$ are pretreatment, is a random sample obeying the following regularity conditions:
@@
These assumptions are on par with prior RD literature, only adding minimal extra regularity to accommodate the covariates $\mathbf{W}_{i}$ for heterogeneity analysis via interactions. The most important substantive condition we require is that $\mathbf{W}_{i}$ are pretreatment, which is formalized in the supplemental appendix (Section SA1.3). Most commonly this holds when $\mathbf{W}_{i}$ is determined before treatment is assigned. Parts (b) and (c) are regularity conditions needed for valid estimation and (robust bias-corrected) inference. Part (c) is a familiar rank condition for least squares regression, adapted to (2.5) and its bias-corrected counterpart. The regularity conditions are slightly stronger than those imposed by Calonico et al. (2019) because of the more flexible regression required for recovery of heterogeneous treatment effects, $\kappa(\mathbf{w})$, compared to the average treatment effect $\tau$ considered in that paper (compare (2.3) to (2.5)).
@@
This assumption requires the potential outcomes to obey a local semilinear-in-$\mathbf{W}$-parameters model, that is, a local partially linear functional coefficient model, which naturally aligns with the estimation procedure in (2.5). It follows from this assumption that the CATE function (at the cutoff $x=c=0$) is linear in $\mathbf{w}$:
@@
From an econometric point of view, Assumption 2 means that estimation of the CATE function remains a one-dimensional nonparametric problem and, further, that the bias can be removed up to the same order as learning the average treatment effect in standard sharp RD settings. This is important for maintaining the empirical tractability of RD analyses. When $\mathbf{W}_{i}$ contains continuous variables, the linearity (in parameters) places a restriction on the “long” regression functions $\mathbb{E}[Y(t)|X_{i}=x,\mathbf{W}_{i}=\mathbf{w}]$, and thus caution must be taken when interpreting the results. Note that the assumption requires linearity in parameters, and so transformations, such as polynomials, can be included in $\mathbf{W}_{i}$ to weaken the implied identifying restrictions. Alternatively, when the covariates are binary, $\mathbf{W}_{i}\in\{0,1\}^{d}$, Assumption 2 is automatically satisfied. Thus, our proposed joint estimation and inference methods are automatically valid for causal heterogeneous RD treatment effects based on discrete covariates (via the construction of orthogonal binary variables for each category), thereby offering a formal causal interpretation for the covariate-interacted local least squares estimates in RD designs.
More generally, Assumption 2 gives a causal interpretation to the commonly used estimation approach in (2.5), but Assumption 1 would be sufficient for all our estimation and inference results if researchers were interested in studying a best mean square approximation of the mean functions $\mathbb{E}[Y(t)|X=c,\mathbf{W}=\mathbf{w}]$—that is, if the parameter of interest were simply the probability limit of the estimated coefficients in (2.5). However, those are often neither useful nor interpretable as heterogeneous causal effects. For more details, see the supplemental appendix.
@@ 3.1 Mean Squared Error and Bandwidth Selection
Estimation and inference rely crucially on selecting a bandwidth $h$ that appropriately localizes to the cutoff $c=0$. This problem is well studied in the methodological literature on RD designs (Arai and Ichimura, 2018; Calonico et al., 2020). The choice of kernel weights is typically less consequential, and typically a linear fit is used for point estimation (see Pei et al., 2021, for more discussion). We develop a bandwidth selection that is optimal in a mean squared error (MSE) sense for estimation of heterogeneous RD treatment effects. That bandwidth yields optimal point estimates of heterogeneous RD treatment effects, making it an appropriate choice for reporting point estimates in empirical applications. However, using such a bandwidth yields invalid inference in general, and thus, in the next section, we develop inference methods using the standard approach of robust bias correction (Calonico et al., 2014).
@@
To characterize the optimal bandwidth for treatment effect estimation, we establish an MSE expansion of the coefficients $\widehat{\bm{\varsigma}}=(\widehat{\theta},\widehat{\bm{\xi}}^{\prime})^{\prime}$ in the local regression (2.5) around their population counterparts. The probability limit of this vector is, under Assumption 2, $\bm{\varsigma}=(\theta(0),\bm{\xi}(0)^{\prime})^{\prime}$. Linear combinations of these yield the CATE function for different groups or contrasts thereof. Let $\bm{\mathsf{s}}$ be the vector that selects or combines the appropriate elements. For example, in the case of a binary $W_{i}$ (so that $d=1$), the base category subgroup can be studied as $\widehat{\theta}=\bm{\mathsf{s}}^{\prime}\widehat{\bm{\varsigma}}$ with $\bm{\mathsf{s}}=(1,0)^{\prime}$, whereas testing group differences would correspond to $\widehat{\bm{\xi}}=\bm{\mathsf{s}}^{\prime}\widehat{\bm{\varsigma}}$ with $\bm{\mathsf{s}}=(0,1)^{\prime}$. In general, setting $\bm{\mathsf{s}}=(1,\mathbf{w})^{\prime}$ for a given value $\mathbf{w}$ yields $\bm{\mathsf{s}}^{\prime}\widehat{\bm{\varsigma}}=\widehat{\theta}+\widehat{\bm{\xi}}^{\prime}\mathbf{w}=\widehat{\kappa}(\mathbf{w})$, recovering the CATE estimator of (2.6).
@@
Here we give the bandwidth selection result. The general MSE expansion is given in the supplemental appendix (Section SA2.6) and is more cumbersome notationally. As it is customary, after some approximations, it depends on the leading variance and squared bias: $\mathsf{Var}[\bm{\mathsf{s}}^{\prime}\widehat{\bm{\varsigma}}]=\frac{1}{nh}\mathsf{V}_{\bm{\mathsf{s}}}$ and $\mathsf{Bias}^{2}[\bm{\mathsf{s}}^{\prime}\widehat{\bm{\varsigma}}]=h^{4}\mathsf{B}_{\bm{\mathsf{s}}}^{2}$, respectively. The rates reflect the specific structure of the estimator in (2.5), that is, a local linear approximation based only on the scalar $X_{i}$. The constants $\mathsf{V}_{\bm{\mathsf{s}}}$ and $\mathsf{B}_{\bm{\mathsf{s}}}$ also capture the specific features of the estimator and target estimand (hence a function $\bm{\mathsf{s}}$). See the supplemental appendix for omitted details. The following theorem summarizes our bandwidth selection result.
@@
$h^{\star}_{\bm{\mathsf{s}}}=\Big{(}\frac{\mathsf{V}_{\bm{\mathsf{s}}}}{4\mathsf{B}_{\bm{\mathsf{s}}}^{2}}\frac{1}{n}\Big{)}^{-1/5},$ (3.1)
@@
While the optimal bandwidth depends on the specific target estimand through the constant terms, its convergence rate is not affected because Equation (2.5) is still a one-dimensional nonparametric (in $X_{i}$) estimation problem. Again, consider the leading case of a binary covariate. If one group is substantially smaller than the other, then all else being equal, the variance will be higher for that group. It could also be that the two groups have conditional expectations with different local curvature, resulting in different bias. Thus, ideally, the bandwidth should be different for each estimand. However, this is not necessarily practical, and furthermore, the specific constant used in bandwidth selection is often less important empirically than ensuring that $h$ is roughly the right magnitude. In the context of heterogeneous effects in RD designs, this perspective supports the common practice of employing the same bandwidth for the estimation of all heterogeneous RD treatment effects of interest. Section 4 provides further discussion in the context of our empirical application.
Our result offers clear practical guidance and justification for several approaches. First, researchers may want to judge if different groups have substantially different bandwidths. If not, it may be fine to conduct all analyses using a common bandwidth. This is easily accommodated by the joint fitting in (2.5). This common bandwidth could be selected as the same bandwidth used for obtaining the average effect in (2.2). One could also use the median of all the relevant bandwidths or, for obtaining robust inference, the smallest of the set. Using different bandwidths for different tasks may yield more accurate inference in some applications.
@@ 3.2 Robust Bias-Corrected Inference
We now state our main results for inference on heterogeneous RD treatment effects. The supplemental appendix gives more general theoretical results (see Section SA2.6), but here we restrict to the case of estimating the joint regression (2.5) using an MSE-optimal bandwidth for point estimation and then conducting inference using robust bias correction. Our main inference procedures rely on a Gaussian distributional approximation for the robust bias corrected versions of $\widehat{\bm{\varsigma}}=(\widehat{\theta},\widehat{\bm{\xi}}^{\prime})^{\prime}$ of the local regression (2.5), along with valid standard errors. This can then be specialized to different use cases depending on which coefficients are combined or contrasted via the continuous mapping theorem.
@@
For inference in RD applications, robust bias correction is a popular way to deal with this problem. In brief, there are two key ingredients to robust bias correction: (i) an estimate of $h^{2}\mathsf{B}$ is subtracted from the point estimate $\widehat{\bm{\varsigma}}$, which amounts to bias correction, and (ii) the standard errors are adjusted to account for the additional estimation error of the bias correction, which amounts to variance correction. Robust bias-correction continues to rely on a Gaussian distributional approximation, but inference is based on the corrected statistics
@@
The robust bias correction method for RD design has been discussed in detail: see Cattaneo et al. (2020), and references therein. See also the supplemental appendix for omitted technical details, including results on heteroskedasdic- and cluster-robust standard errors (Cameron and Miller, 2015; MacKinnon et al., 2023). The method is known to have excellent theoretical and numerical properties (Calonico et al., 2018, 2022), and has been shown to yield correct inference in extensive validation empirical studies (Hyytinen et al., 2018; De Magalhães et al., 2025).
@@ 4 Application
To illustrate our methodological results we revisit the analysis by Akhtari et al. (2022), which examines the impact of political turnover in Brazilian mayoral elections on local public service provision using an RD design based on close elections. Klašnja and Titiunik (2017) first studied this data and RD design, and documented large negative RD effects of party incumbency based on Brazilian mayoral elections, comparing municipalities where the incumbent party barely loses (resulting in political turnover) to those where the incumbent party barely wins (no turnover): a bare victory results in a large reduction in the probability of victory in the following election, which they attribute to deficient accountability resulting from weak parties and term limits.
@@
The analysis of Akhtari et al. (2022) highlights two key methodological difficulties researchers have faced when studying heterogeneous treatment effects in RD without methodological and theoretical guidance in the literature, as discussed above. First, Akhtari et al. (2022) use an optimal data-driven bandwidth selection method (Calonico et al., 2019) to localize at the cutoff. However, both estimation and inference rely solely on local linear fitting. Although estimation using linear regression (specifically, their implementation of (2.3)) is correct, valid inference requires robust bias correction, as standard errors from the linear fit are invalid when paired with the optimal bandwidth. In our replication, we revisit their heterogeneity analysis using robust bias-corrected inference. Our results, presented in Table 1, report confidence intervals instead of standard errors, following best practices in RD analysis.
@@
The second key aspect is that the heterogeneity variable of interest, income, is continuous. Akhtari et al. (2022) binarize income and compare municipalities above and below the median income level. We first replicate this exercise using our methods. The results are shown in the top two rows of Panel (b), Table 1. Our point estimates are nearly identical to theirs (0.373 and 0.122 compared 0.389 and 0.126 in their Table A.21), where the differences are due to slight bandwidth variations. Next, we incorporate valid uncertainty measures using robust bias correction, revealing that the treatment effect is statistically significant only for low (below-median) income municipalities. We also formally test the null hypothesis that the treatment effect is the same across income groups, finding strong evidence against it: our robust inference yields a 95% confidence interval of [-0.422, -0.078] for the difference, which excludes zero thus revealing statistically significant heterogeneity. We also extend this discretized analysis by studying treatment effects by income quartiles and deciles, rather than a binary median split. This refinement adds nuance to the original study, showing that variation in treatment effects is more pronounced across the median than within the upper or lower groups. These results further strengthen the original findings.
@@
Finally, we analyze income directly, as a continuous variable. The specification remains (2.5), but its interpretation is now governed by Assumption 2: with an intercept and slope coefficient. In Table 1, Panel (c) reports estimated intercept and slope based on a linear-in-income model with $\mathbf{W}_{i}=\mathtt{income}_{i}$, while Panel (d) considers a quadratic-in-income model with $\mathbf{W}_{i}=(\mathtt{income}_{i},\mathtt{income}_{i}^{2})$. Comparing the discretized heterogeneity with this linear and quadratic models, as shown graphically in Figure 1, reveals that the two continuous models effectively capture the overall heterogeneity in this application. Notably, the continuous specifications highlight differences in income distribution within each group, particularly emphasizing that the top quartile has a much wider range than the others—an important finding for policy implications.
@@
Ultimately, the decision of how to best capture heterogeneity is up to the researcher. The discretized and continuous versions have their strengths and limitations. The continuous approach, particularly when compared to a fine discretization, may yield substantial precision improvements, but requires the structure of Assumption 2. The choice ultimately depends on the specific application and the researcher’s objectives, but in any case, it is important to have rigorous estimation and inference methods for both.

### v3

@@ 2 Setup
Unlike Equation (2.3), our proposed fit (2.5) does not yield an estimator of the average treatment effect (2.1). This is implicit in Theorem 1 below, and was already shown by Calonico et al. (2019). There are two potential remedies to this, neither of which we recommended in practice. First, the covariates can be demeaned on each side of the cutoff. This would follow the logic for linear regression adjustments in randomized experiments, but in this case, the demeaning would amount to Nadaraya–Watson estimation at a boundary point with $\mathbf{W}_{i}$ as the outcome, thereby hampering the entire RD estimation due to the presence of smoothing bias. Second, one could average the fitted CATEs, but this again essentially amounts to Nadaraya–Watson estimation at a boundary point, now using $\widehat{\kappa}(\mathbf{w})$ as the outcome. In both cases, the convergence rate will be slow because the misspecification bias is of order $h$, and is thus impractical for inference (e.g., severe undersmoothing would be needed). Therefore, in applications, researchers should only use our proposed methods for heterogeneity analysis and rely on the established methods underlying (2.2) or (2.3) to estimate the average RD treatment effect.
Finally, a related motivation for the specification (2.7) is to control for panel or group structures. It is common in some RD applications to include fixed effects for individuals, groups, or time periods. Identification, estimation, and inference then follow from the same arguments presented in the upcoming sections. In this way, our proposed RD methodology, although developed for the cross-sectional setting, can be easily applied to panel data, provided weighted linear least squares methods with additive or interactive fixed effects are used.
@@ 3 Main Results
We turn to the theoretical properties of the regression fit in (2.5), and particularly the CATE estimator (2.6). We provide a complete set of practical tools for identification, estimation, and inference. The following standard regularity conditions are imposed. Recall that the cutoff is $c=0$ without loss of generality, and let $x_{\mathtt{L}}<0<x_{\mathtt{U}}$. All omitted details are given in the supplemental appendix (see Section SA1.3).
@@
The data $(Y_{i},X_{i},\mathbf{W}_{i}^{\prime}),i=1,\ldots,n$, where $T_{i}=\mathbbm{1}\{X_{i}\geq 0\}$ and $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$, are a random sample obeying the following:
@@
These assumptions are on par with prior RD literature, only adding minimal extra regularity to accommodate the covariates $\mathbf{W}_{i}$ for heterogeneity analysis via interactions. The most important substantive condition we require is that $\mathbf{W}_{i}$ are pretreatment. In particular, continuity of $\mathbb{E}[\mathbf{W}_{i}|X_{i}=x]$ means there is no RD treatment effect on the covariates, while the same is required for higher moments for the asymptotic analyses. Most commonly, this holds when $\mathbf{W}_{i}$ is determined before treatment is assigned. Parts (b) and (c) are regularity conditions needed for valid estimation and (robust bias-corrected) inference. Part (c) is a familiar rank condition for least squares regression, adapted to (2.5) and its bias-corrected counterpart. The regularity conditions are slightly stronger than those imposed by Calonico et al. (2019) because of the more flexible regression required for recovery of heterogeneous treatment effects, $\kappa(\mathbf{w})$, compared to the average treatment effect $\tau$ considered in that paper (compare (2.3) to (2.5)). Part (d) is a standard assumption when employing kernel-weighted nonparametric regressions.
@@
This assumption requires the potential outcomes to obey a local semilinear-in-$\mathbf{W}$-parameters model, that is, a local linear functional coefficient model, which naturally aligns with the estimation procedure in (2.5). It follows from this assumption that the CATE function (at the cutoff $x=c=0$) is linear in $\mathbf{w}$:
@@
From an econometric point of view, Assumption 2 means that estimation of the CATE function remains a one-dimensional nonparametric problem and, further, that the bias can be removed up to the same order as learning the average treatment effect in standard sharp RD settings. This is important for maintaining the empirical tractability of RD analyses. When $\mathbf{W}_{i}$ contains continuous variables, the linearity (in parameters) places a restriction on the “long” regression functions $\mathbb{E}[Y_{i}(t)|X_{i}=x,\mathbf{W}_{i}=\mathbf{w}]$, and thus caution must be taken when interpreting the results. Note that the assumption requires linearity in parameters, and so transformations, such as polynomials, can be included in $\mathbf{W}_{i}$ to weaken the implied identifying restrictions. Alternatively, when the covariates are binary, $\mathbf{W}_{i}\in\{0,1\}^{d}$, Assumption 2 is automatically satisfied. Thus, our proposed joint estimation and inference methods are automatically valid for causal heterogeneous RD treatment effects based on discrete covariates (via the construction of orthogonal binary variables for each category), thereby offering a formal causal interpretation for the covariate-interacted local least squares estimates in RD designs.
More generally, Assumption 2 gives a causal interpretation to the commonly used estimation approach in (2.5), but Assumption 1 would be sufficient for all our estimation and inference results if researchers were interested in studying a best mean square approximation of the mean functions $\mathbb{E}[Y_{i}(t)|X_{i}=c,\mathbf{W}_{i}=\mathbf{w}]$—that is, if the parameter of interest were simply the probability limit of the estimated coefficients in (2.5). However, those are often neither useful nor interpretable as heterogeneous causal effects. For more details, see the supplemental appendix.
@@ 3.1 Mean Squared Error and Bandwidth Selection
Estimation and inference rely crucially on selecting a bandwidth $h$ that appropriately localizes to the cutoff $c=0$. This problem is well studied in the methodological literature on RD designs (Arai and Ichimura, 2018; Calonico et al., 2020). The choice of kernel weights is typically less consequential, and typically a linear fit is used for point estimation (see Pei et al., 2021, for more discussion). We develop a bandwidth selection method that is optimal in a mean squared error (MSE) sense for estimation of heterogeneous RD treatment effects. Such bandwidth yields optimal point estimates of heterogeneous RD treatment effects, making it an appropriate choice for reporting point estimates in empirical applications. However, using such a bandwidth yields invalid inference in general, and thus, in the next section, we develop inference methods using the standard approach of robust bias correction (Calonico et al., 2014).
@@
To characterize the optimal bandwidth for treatment effect estimation, we establish an MSE expansion of the coefficients $\widehat{\bm{\varsigma}}=(\widehat{\theta},\widehat{\bm{\xi}}^{\prime})^{\prime}$ in the local regression (2.5) around their population counterparts. The probability limit of this vector is, under Assumption 2, $\bm{\varsigma}=(\theta(0),\bm{\xi}(0)^{\prime})^{\prime}$. Linear combinations of elements of $\bm{\varsigma}$ yield the CATE function for different groups or contrasts. Let $\bm{\mathsf{s}}$ be the conformable vector that selects or combines the appropriate elements. For example, in the case of a binary $W_{i}$ ($d=1$), the base category subgroup can be studied as $\widehat{\theta}=\bm{\mathsf{s}}^{\prime}\widehat{\bm{\varsigma}}$ with $\bm{\mathsf{s}}=(1,0)^{\prime}$, whereas testing group differences would correspond to $\widehat{\bm{\xi}}=\bm{\mathsf{s}}^{\prime}\widehat{\bm{\varsigma}}$ with $\bm{\mathsf{s}}=(0,1)^{\prime}$. In general, setting $\bm{\mathsf{s}}=(1,\mathbf{w})^{\prime}$ for a given $\mathbf{w}$ yields $\bm{\mathsf{s}}^{\prime}\widehat{\bm{\varsigma}}=\widehat{\theta}+\widehat{\bm{\xi}}^{\prime}\mathbf{w}=\widehat{\kappa}(\mathbf{w})$, recovering the CATE estimator of (2.6).
@@
We now discuss the bandwidth selection result. The general MSE expansion is more cumbersome notationally; we provide it in the supplemental appendix (Section SA2.6). As is customary, after some approximations, it depends on the leading variance and squared bias: $\mathsf{Var}[\bm{\mathsf{s}}^{\prime}\widehat{\bm{\varsigma}}]=\frac{1}{nh}\mathsf{V}_{\bm{\mathsf{s}}}$ and $\mathsf{Bias}^{2}[\bm{\mathsf{s}}^{\prime}\widehat{\bm{\varsigma}}]=h^{4}\mathsf{B}_{\bm{\mathsf{s}}}^{2}$, respectively. The rates reflect the specific structure of the estimator in (2.5), a local linear approximation based only on the scalar $X_{i}$. The constants $\mathsf{V}_{\bm{\mathsf{s}}}$ and $\mathsf{B}_{\bm{\mathsf{s}}}$ also capture the specific features of the estimator and target estimand (hence they are functions of $\bm{\mathsf{s}}$). The following theorem summarizes our bandwidth selection result.
@@
$h^{\star}_{\bm{\mathsf{s}}}=\Big{(}\frac{\mathsf{V}_{\bm{\mathsf{s}}}}{4\mathsf{B}_{\bm{\mathsf{s}}}^{2}}\frac{1}{n}\Big{)}^{1/5},$ (3.1)
@@
While the optimal bandwidth depends on the specific target estimand through the constant terms, its convergence rate is not affected because equation (2.5) is still a one-dimensional nonparametric (in $X_{i}$) estimation problem. Again, consider the leading case of a binary covariate. If one group is substantially smaller than the other, then, all else equal, the variance will be higher for that group. It could also be that the two groups have conditional expectations with different local curvature, resulting in different biases. Thus, ideally, the bandwidth should be different for each estimand. However, this is not necessarily practical, and the specific constant used in bandwidth selection is often less important empirically than ensuring that $h$ is roughly of the right magnitude. In the context of heterogeneity analysis in RD designs, this perspective supports the common practice of employing the same bandwidth for the estimation of all heterogeneous RD treatment effects. Section 4 provides further discussion in the context of our empirical application.
Our result offers clear practical guidance and justification for several approaches. Researchers may decide that different groups should have substantially different bandwidths. Alternatively, they may prefer to conduct all analyses using a common bandwidth. This is easily accommodated by the joint fitting in (2.5). This common bandwidth could be selected as the same bandwidth used for the average effect in (2.2). One could also use the median of all the relevant bandwidths or, for robust inference, the smallest of the set. Using different bandwidths for different tasks may yield more accurate inference in some applications.
@@ 3.2 Robust Bias-Corrected Inference
We now state our main results for inference on heterogeneous RD treatment effects. We consider the case of estimating the joint regression (2.5) using an MSE-optimal bandwidth for point estimation and then conducting inference using robust bias correction—the supplemental appendix gives more general theoretical results (see Section SA2.6). Our main inference procedures rely on a Gaussian distributional approximation for the robust bias corrected versions of $\widehat{\bm{\varsigma}}=(\widehat{\theta},\widehat{\bm{\xi}}^{\prime})^{\prime}$ of the local regression (2.5), along with valid standard errors. This can then be specialized to different use cases depending on which coefficients are combined or contrasted via the continuous mapping theorem.
@@
For inference in RD applications, robust bias correction is a popular way to deal with this problem. In brief, there are two key ingredients to robust bias correction: (i) an estimate of $h^{2}\mathsf{B}$ is subtracted from the point estimate $\widehat{\bm{\varsigma}}$, which amounts to bias correction, and (ii) the standard errors are adjusted to account for the additional estimation error of the bias correction step, which amounts to variance correction. Robust bias-correction continues to rely on a Gaussian distributional approximation, but inference is based on the corrected statistics
@@
The robust bias correction method for RD design has been discussed in detail, see Cattaneo et al. (2020), and references therein. See also the supplemental appendix for omitted technical details, including results on heteroskedastic- and cluster-robust standard errors (Cameron and Miller, 2015; MacKinnon et al., 2023). Robust bias correction is known to have excellent theoretical and numerical properties (Calonico et al., 2018, 2022), and has been shown to yield correct inference in extensive validation empirical studies (Hyytinen et al., 2018; De Magalhães et al., 2025).
@@ 4 Application
To illustrate our methodological results, we revisit the analysis by Akhtari et al. (2022), which examines the impact of political turnover in Brazilian mayoral elections on local public service provision using an RD design based on close elections. Klašnja and Titiunik (2017) first studied these data and RD design, and documented large negative RD effects of party incumbency based on Brazilian mayoral elections, comparing municipalities where the incumbent party barely loses (resulting in political turnover) to those where the incumbent party barely wins (no turnover): a bare victory results in a large reduction in the probability of victory in the following election, which they attribute to deficient accountability resulting from weak parties and term limits.
@@
The analysis of Akhtari et al. (2022) highlights two key methodological difficulties researchers have faced when studying heterogeneous treatment effects in RD without methodological and theoretical guidance in the literature, as discussed above. First, Akhtari et al. (2022) use an optimal data-driven bandwidth selection method (Calonico et al., 2019) to localize at the cutoff. However, both estimation and inference rely solely on local linear fitting. Although estimation using linear regression (specifically, their implementation of (2.3)) is correct, valid inference requires robust bias correction, as standard errors from the linear fit are invalid when paired with the optimal bandwidth. In our replication, we revisit their heterogeneity analysis using robust bias-corrected inference. Our results, presented in Table 1, report confidence intervals instead of standard errors, following best practices for RD analysis.
@@
The second key aspect is that the heterogeneity variable of interest, income, is continuous. Akhtari et al. (2022) binarize it to compare municipalities above and below the median income level. We first replicate this exercise using our methods. The results are shown in the top two rows of Panel (b), Table 1. Our point estimates are nearly identical to theirs (0.373 and 0.122 compared 0.389 and 0.126 in their Table A.21), where the differences are due to slight bandwidth variations. Next, we incorporate valid uncertainty measures using robust bias correction, revealing that the treatment effect is statistically significant only for low (below-median) income municipalities. We also formally test the null hypothesis that the treatment effect is the same across income groups, finding strong evidence against it: our robust inference yields a 95% confidence interval of [-0.422, -0.078] for the difference, which excludes zero and reveals statistically significant heterogeneity. We also extend this discretized analysis by studying treatment effects by income quartiles and deciles, rather than a binary median split. This refinement adds nuance to the original study, showing that variation in treatment effects is more pronounced across the median than within the upper or lower groups. These results further strengthen the original findings.
@@
Finally, we analyze income directly, as a continuous variable. The specification remains (2.5), but its interpretation is now governed by Assumption 2 with an intercept and slope coefficient. In Table 1, Panel (c) reports estimated intercept and slope based on a linear-in-income model with $\mathbf{W}_{i}=\mathtt{income}_{i}$, while Panel (d) considers a quadratic-in-income model with $\mathbf{W}_{i}=(\mathtt{income}_{i},\mathtt{income}_{i}^{2})$. Comparing the discretized heterogeneity with these linear and quadratic models, as shown graphically in Figure 1, reveals that the two continuous models effectively capture the overall heterogeneity in this application. Notably, the continuous specifications highlight differences in income distribution within each group, particularly emphasizing that the top quartile has a much wider range than the others—an important finding with policy implications.
@@
Ultimately, the decision of how to best capture heterogeneity is up to the researcher. The discretized and continuous versions have their strengths and limitations. The continuous approach, particularly when compared to a fine discretization, may yield substantial precision improvements, but requires imposing the structure of Assumption 2. The choice ultimately depends on the specific application and the researcher’s objectives. Our framework provides rigorous estimation and inference methods for both cases.